OpenStudy (anonymous):

When a set 4 integers are summed together in pairs, the resulting sums are 10, 13, 16, 19, 22, and 25. What is their product?

OpenStudy (raden):

let that number are a, b, c, and d. the equations are : a+b=10 a+c=13 a+d=16 b+c=19 b+d=22 c+d=25

OpenStudy (raden):

add the 1st equation with the 2nd, get a+b=10 a+c=13 ------- + 2a + b+c = 23 in the 4th equation says that b+c = 19, so 2a + b+c = 23 2a + 19 = 23 2a = 23 - 19 2a = 4 a = 4/2 = 2 i think the rest be easy for you now to get the values of b,c, and d

OpenStudy (raden):

then multiply of them a * b * c * d = ...